Hypothesis Testing, P-values: For the given alternate hypothesis (H1) and test statistic (z), determine the P-value of the test statistic. Use the answer found in the z-table or round to 4 decimal places.

(a)    H1: p < 0.75
The test statistic is

z = −2.83.


P-value =

(b)    H1: p > 0.85
The test statistic is

z = 1.84.


P-value =

(c)    H1: p ≠ 0.15
The test statistic is

z = 1.85.


P-value =
Answer #1

solution

(A)this is left tailed test

P(z < −2.83 ) = 0.0023

P-value =0.0023

(B)this is right tailed test

P(z >1.84 ) = 1 - P(z < 1.84) = 1-0.9671=0.0329

P-value = 0.0329

(C)this is two tailed test

P(z >1.85 ) = 1 - P(z < 1.85) = 1-0.9678=0.0322=0.0644

P-value = 0.0644

( note = using z table )